OpenStudy (anonymous):

Solve for t using logarithms with base a. \[3a^{4t}=10\]

OpenStudy (anonymous):

Use this : \[Loga^b = b.Loga\]

OpenStudy (anonymous):

Okay, I'm going to put all my steps... sorry...

OpenStudy (anonymous):

\[log_a(3a^{4t}=log_a(10))\] \[log_a(3)+log_a(a^{4t})=log_a(10)\] \[4t=log_a(10)-log_a(3)\] \[4t=log_a(\frac{10}{3})\] \[t=\frac{1}{4}(log_a(\frac{10}{3}))\]

OpenStudy (anonymous):

What is the answer???

OpenStudy (anonymous):

The last line... or do I have to multiply it out...?

OpenStudy (anonymous):

How can you multiply it?? I think it will go into the power..

OpenStudy (anonymous):

Yeah but either way I still got the wrong answer...

OpenStudy (anonymous):

I am asking for what is the correct answer??

OpenStudy (anonymous):

Oh it's an even question but wolfram has the answer: \[t=\frac{log(\frac{10}{3})}{4 log (a)}\] But is there more than one way to do this because another website also gives has the answer: \[t=\frac{log_a(10)}{4log_a(3a)}\]

OpenStudy (anonymous):

Second one I have got...

OpenStudy (anonymous):

Is there more than one to do this... or anyway you do it your supposed to come up with the same answer?

OpenStudy (anonymous):

You have made a mistake there.. are you able to point out or not???

OpenStudy (anonymous):

No am not...

OpenStudy (anonymous):

Why you taken log to the base a?? Well, it is not necessary.. You can take it as log to the base 10 also..

OpenStudy (anonymous):

By taking log to the base 10. you will get: \[Log(a^{4t}) = 4t.Log(a)\] Use this you will get your answer..

OpenStudy (anonymous):

I show both the methods one by one just wait:

OpenStudy (anonymous):

Well, one... those are the instructions and two...Because I thought that you have to use the cancellation law if you have an exponent a^x to you raise it to the log_a to find x or make it equal to whatever, x, is in the exponent.

OpenStudy (anonymous):

Okay :).

OpenStudy (anonymous):

But the instructions state using logarithms with base a

OpenStudy (anonymous):

First Method: Take Log both the sides: \[Log(3^{4t}) = Log(10)\] \[Log(3) + Log(a^{4t}) = Log(10)\] \[4t.Log(a) = Log(10) - Log(3)\] \[4t.Log(a) = Log(\frac{10}{3})\] \[t = \frac{Log(\frac{10}{3})}{4.Log(a)}\]

OpenStudy (anonymous):

Yes, second method is using that instruction..Just wait I will explain..

OpenStudy (anonymous):

Have you left out the first line of the by accident \[log_a(3a^{4t})\] or...

OpenStudy (anonymous):

Take Log(to the base a both sides), \[Log_a(3^{4t}) = Log_a(10)\] \[4t.Log_a(3a) = Log_a(10)\] So, from here evaluate t: \[t = \frac{Log_a(10)}{4.Log_a(3a)}\] Here is your second answer.. You can do more here if you want to.. Just carry the 4 to the power of (3a), you will get 81a^4.. Getting or not??

OpenStudy (anonymous):

No... for the first one... can't I use the cancellation law that state \[log_a(a^x)=x\] to get rid of the a...

OpenStudy (anonymous):

Sorry, in first line it will come 3a^(4t) and not 3^(4t).. You can go with your method also there is nothing wrong in that...

OpenStudy (anonymous):

But... I thought regardless of the method am suppose to get the same answer?

OpenStudy (anonymous):

See, the answer provided by wolfram and the other website is also different.. Your answer can also be different.. What is the problem in that???

OpenStudy (anonymous):

Sigh, but their answers are very similar...

OpenStudy (anonymous):

See, what more can you do is: \[4.Log_a(3a) = 4.Log_a(3) + 4.Log_a(a) = 4(Log_a(3) + 1)\]

OpenStudy (anonymous):

:\

OpenStudy (anonymous):

nothing wrong with what you have in your first post.... \(\large t=\frac{1}{4}(log_a(\frac{10}{3})) \)

OpenStudy (anonymous):

but you can also utilize the properties of logs: \(\large t=\frac{1}{4}(log_a(\frac{10}{3})) \) \(\large t=\frac{1}{4}(log_a10-log_a3) \) \(\large t=\frac{1}{4}log_a10-\frac{1}{4}log_a3 \)

OpenStudy (anonymous):

Or it can be: \[t = Log_a(10)^{1/4} - Log_a(3)^{1/4}\]

OpenStudy (anonymous):

yes^^^ :)

OpenStudy (anonymous):

How would I convert that back to the original?

OpenStudy (anonymous):

just work backwards from the formula/properties

OpenStudy (anonymous):

Okay, I'll put it in another question. Thanks so much for all your help I really do appreciate it everyone!!

OpenStudy (anonymous):

but also, just to be clear with what waterineyes wrote, the expression is: \[\huge t=Log_a(10^\frac{1}{4})-Log_a(3^\frac{1}{4}) \]

OpenStudy (anonymous):

@dpaInc how you write this in so large form? can you please tell???

OpenStudy (anonymous):

instead of "\large", use "\huge" in your latex...:)